An economy is defined by
C= 400+0.4(Y-T)
IP= 200
NX= 40
T=0
Y*=1500
Calculate autonomous expenditure, multiplier, short run equilibrium and output gap.
1
Expert's answer
2019-11-27T09:55:44-0500

Given; C = (consumption)"C = 400 + 0.4 (Y-T)"

IP = 200(investment)

NX = 40 (net exports)

T = 0(taxes)

Y* = 1500(potential output)

To get the actual output we use the formula;"Y = C+I+G+N" .

Substituting we get: Y=400+0.4Y-0.4T+I+G+NX

0.6Y=400-0.4T+I+G+NX; Divide everything with 0.6 to get value of Y.

Y=667.67+0.67T+1.67I+1.67G+1.67NX

Now replace with the values of T, I, G and NX to get;

Y=667.67+0.67+334+1.67+66.8

Y=1069.81

Output gap is given by;"Y-Y*" . Which is 1069.81-1500 =-430.19

Tax multiplier is the value of T in the above equation which is =0.67

Autonomous expenditure, "AE=C"  which is (400+0.4*1069.81) =827.92